OpenStudy (anonymous):

solve by completing te square 5x^2=7x-4

OpenStudy (anonymous):

When they say completing the square what comes to mind?

OpenStudy (anonymous):

taking half of the coefficient and squaring it

OpenStudy (anonymous):

coefficient of the x term

OpenStudy (anonymous):

Actually, yes. You want to get everything with an x to one side: 5x^2-7x=4 The bad part about completing the square is that the x^2 has to have a coefficient of 1 so now you "get" to divide it all by 5: x^2-7x/5=4/5 Now add half the x's coeffiecent to both sides.